OpenStudy (anonymous):

Please factor X^2=-2x+33 show your steps! medal and fan reward

OpenStudy (anonymous):

are you asked to factor this, or find the zeros?

OpenStudy (anonymous):

facotr

OpenStudy (anonymous):

Well, in a sense, finding the zeros and factoring accomplish the same things, in order to find the factor for this, you need to use the Quadratic Equation, recall what that it\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]where your a,b, and c all come from the general form of a polynomial that looks like this\[ax^2+bx+c=0\]so you need to first move all terms to one side, then plug your a,b, and c into the quadratic equation, and that will give you your factors

OpenStudy (anonymous):

can u show the other steps? @Embryo

OpenStudy (anonymous):

alrighty, so first thing to do is to move everything to one side, after adding 2x to both sides, and subtracting 33 from both sides, you get this expression\[x^2+2x-33=0\]now that it's in this form, you can see that your values for a,b, and c are, a=1,b=2, and c=-33 once you plug these into the Quadratic Equation you'll get your two answers, and you'll have your factors\[\frac{ -2 \pm \sqrt{2^2-4(1)(-33)} }{ 2(1) }=x\]